OpenStudy (anonymous):

Which equation is not equivalent to the formula m = ca? c=m/a a=c/m a=m/c m = ac

OpenStudy (anonymous):

you got answer choices ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

m = ca c = m/a -- equivalent a = c/m -- not equivalent a = m/c -- equivalent m = ac -- equivalent

OpenStudy (anonymous):

thank you :)